OpenStudy (anonymous):

3 = log8 + 3logx. Solve for x.

OpenStudy (anonymous):

i dont no what it is

OpenStudy (anonymous):

Me either

OpenStudy (anonymous):

\(\large 3=log8+3logx \) \(\large 3=log8+log(x^3) \) \(\large 3=log(8\cdot x^3) \) now write that last line in exponential form... can you do that?

OpenStudy (anonymous):

Would it be 8^x3=3?

OpenStudy (anonymous):

no... that problem is the common log right? base 10?

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

so where's the 10 in your equation?

OpenStudy (anonymous):

It's not shown. I thought there would be one but that's all it had :/

OpenStudy (anonymous):

\[\huge log_bM=x \Rightarrow b^x=M \]

OpenStudy (anonymous):

your base is 10, and your exponent is the 3. the M is whatever is in the parenthesis...

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

I think I got how to do it

OpenStudy (anonymous):

so your equation in exponential form should be: \(\large 10^3=8x^3 \)

OpenStudy (anonymous):

ok.... good luck to you...